# Three-view Focal Length Recovery From Homographies

### v1

@@ 1 Introduction
Estimating relative camera motion from multiple views using point correspondences is a classical problem in computer vision. Efficient solutions for various camera configurations have been well-studied in the literature [37, 19, 27, 43, 33, 4, 18, 20, 21, 38]. For example, with two fully calibrated cameras, the relative camera pose can be efficiently determined using 5-point algorithms [37, 19, 27]. In contrast to the two view case, the calibrated three-view relative pose problem is much more challenging. This problem can be solved using four triplets of point correspondences [38, 39], [^1] Note, that configuration of four points in three views generates an over-constrained problem. In this case, we have 12 constraint for 11 degrees of freedom (DOF). A minimal solution would need to drop one constraint, e.g., by considering only a line passing through one of the points in the third view [24] or by considering a “half” point correspondence. however it results in a very complex system of polynomial equations. Thus, the existing solutions [38, 24] to this problem are only approximate and can often fail, i.e., the returned solution can be arbitrarily far from the geometrically correct solution.
@@
An important scenario arises when the focal length is the only unknown in the camera intrinsic parameters, referred to as the partially calibrated case. This setup is practical since most modern cameras have zero skew and a centered principal point. When two cameras share an equal yet unknown focal length, their relative motion and the common focal length can be estimated using six point correspondences [43, 19, 27, 29]. Similarly, if one camera is fully calibrated and the other has an unknown focal length, six point correspondences are also required [5, 27, 29]. For two cameras with different, unknown focal lengths, at least seven point correspondences are necessary to recover both the relative motion and the focal lengths [18, 3].
Three-view focal length problems are significantly more complex, often involving polynomial systems with hundreds of solutions. For example, when the three cameras share the same focal length, there can be up to 668 possible solutions [7, 13, 10]. Such systems can be solved using homotopy continuation methods [41, 22]. The running times of the CPU variants of the solvers for four points in three cameras with unknown shared focal length range from $250ms$ to $1456ms$. Efficient GPU implementations are much faster; however, with runtimes $16.7ms$ to $154ms$ they are still too slow for practical applications.
Scenes with planar surfaces, such as floors, walls, doors, facades, and other common structures, are prevalent in man-made environments. When points are coplanar, homography-based algorithms require four point correspondences [20] to estimate the relative pose of two cameras. Unfortunately, any attempt to recover intrinsic parameters from two views of a planar surface (using only point correspondences without additional priors) is futile, as stated in the following theorem [36, 37]: ”For any choice of intrinsic parameters, any homography can be realized between two views by positioning the views and a plane in a particular way.” However, it has been shown that focal lengths can be recovered from three-view homographies [35, 23], i.e. three cameras observing a planar surface (see illustration in Figure 1).
@@
  • This allows us to develop solvers using efficient Sturm sequences or hidden variable technique. The proposed solvers are significantly more efficient than the existing solvers to this problem [23]. Moreover, in extensive synthetic and real experiments, we show an improvement in accuracy over the state-of-the-art two-view solvers.
  • We propose a new dataset consisting of 6 scenes (four indoor and 2 outdoor) containing 1870 images captured with 14 different cameras. Ground truth focal lengths are estimated using the standard calibration method [46]. We will make both the dataset and the code publicly available.
@@ 2 Related Work
Homography estimation is a well-studied problem in the literature and can be solved using the 4-point algorithm, which involves solving a system of eight homogeneous linear equations [20]. However, with only point correspondences, two-view homography does not provide additional constraints on the focal length, making the two-view homography not suitable for self-calibration. To recover the intrinsic parameters from the two-view homography, we need additional priors. In [4], under pure rotation assumption, the authors propose two minimal solutions for the two-view homography-based focal length estimation with one or two focal length parameters. In [15], the authors use sensor-fusion, i.e., combining camera with gravity from IMU (Inertial measurement unit), to reduce the DOF of the motion parameters. In general, the gravity prior can provide up to two constraints on the focal length.
Focal length recovery from homographies in more views was first discussed in [35]. Malis et al. compute homographies among multiple views, iteratively testing various focal lengths to find the one that minimizes a cost function. However, this method is time-consuming and sensitive to noise.
The work most closely related to ours is by Heikkilä [23], where the constraints on the focal length and the normal vector are derived, and non-iterative solutions to the problem of recovering focal lengths from three-view homographies is proposed. For three cameras with the same unknown focal length,  [23] solves three polynomials in three unknowns. However, these polynomials have very high degrees, making them challenging to solve efficiently. Heikkilä first applies SVD to a $82\times 82$ matrix, followed by Gaussian-Jordan elimination of a $82\times 164$ matrix. Finally, the solver finds the solutions by computing the eigenvalues of a $82\times 82$ matrix. For three cameras with two different focal lengths, Heikkilä’s method needs to solve four polynomials in four unknowns by performing Gaussian-Jordan elimination of a $176\times 352$ matrix and computing the eigenvalues of a $176\times 176$ matrix.
In contrast to complex solutions presented in [23], we propose very efficient solutions to the problem of recovering focal lengths from three-view homographies. For cameras with the same focal length, our novel solver only requires computing the roots of a univariate polynomial of degree 9, which can be efficiently solved using Sturm sequences. For cameras with two different focal lengths, we compute solutions from the eigenvalues of a $18\times 18$ matrix.
@@
Problems Proposed Solvers View 1 View 2 View 3 Ours Heikkilä [23] Ours Heikkilä [23] Ours Heikkilä [23]
@@
Case IV $\mathbf{H}_{f\rho}$ Known $f$ $\rho$ 9 - $12\times 12$ 106 -
@@
For general scenes, focal lenghts can be recovered together with the epipolar geometry. For two cameras with equal unknown focal lengths, the minimal 6-point solvers [43, 5, 19, 27, 29] use rank-2 and trace constraints on the essential matrix. The 5-point solver [44] solves the plane+parallax scenario for cameras with unknown focal length using four coplanar points and one off-plane point. This solver can be used within the DEGENSAC framework [9] to complement the six point solvers in order to improve performance for scenes that are not completely planar, but contain a dominant plane. For the single unknown focal length problem, the two-view geometry can be solved using the single-side 6-point solver [5, 29]. For the relative pose problem with different and unknown focal lengths, the solution computes the fundamental matrix using the 7-point solver [20], followed by the extraction of the focal lengths from this matrix [3, 25].
As discussed above, three-view focal length problems for general scenes are significantly more complex and their homotopy continuation solutions [7, 13, 10] are impractical.
@@ 3 Problem Statement
The points are observed by three cameras such that $\mathbf{X}_{i}$ is projected to a 2D point $\mathbf{m}_{i,j}$ by the $j$-th camera $\mathbf{K}_{j}[\mathbf{R}_{j}\ |\ \mathbf{t}_{j}]$. In many practical scenarios, it is often reasonable to assume that the cameras have square-shaped pixels, and the principal point coincides with the image center [19]. This leaves only the focal point unknown and thus $\mathbf{K}_{j}={\rm diag}(f_{j},f_{j},1)$.
@@ 4 Our Approach
$\begin{split}\mathbf{H}_{j}\sim\mathbf{K}_{j}^{-1}\mathbf{G}_{j}\mathbf{K}_{1}.\\ \end{split}$ (7)
Thus for known $\mathbf{G}_{j}$, $\mathbf{H}_{j}$ are polynomial matrices in the focal length parameters. To solve for the focal lengths, we need to find the constraints on $\mathbf{H}_{j}$. Based on the formulation of the Euclidean homography (4), we know that [20]
@@
Thus the essential matrices derived from different $\mathbf{H}_{j}^{\top}$ are related by the same normal vector $\mathbf{n}$ in the reference camera coordinate. As shown in [42, 37], a valid essential matrix should satisfy the singular and trace constraints
@@
Substituting (7) into the trace constraints (11) gives us nine equations per homography, i.e., 18 equations for two homography matrices $\mathbf{H}_{2}$ and $\mathbf{H}_{3}$. Only two from the nine equations are algebraically independent, i.e., one homography matrix provides two constraints on the focal lengths and the normal vector parameters. For the case of equal unknown focal length, we have only 3 unknowns $(f,n_{x},n_{y})$ (since (11) is homogenous in $\mathbf{n}$, we can let $n_{z}=1$) and thus an over-constrained problem. The problem can be solved using only one of the nine trace constraint equations for $\tilde{\mathbf{E}}_{2}$. These equations can be solved using the Gröbner basis method [30, 32] The final solver performs Gauss-Jordan elimination of a $118\times 141$ matrix and extracts solutions from the eigenvectors of a $23\times 23$ matrix. However, this solver is not efficient enough for real-time applications.
@@
$\begin{split}[\mathbf{n}]_{\times}\mathbf{H}_{j}^{\top}\mathbf{H}_{j}[\mathbf{n}]_{\times}^{\top}&=[\mathbf{n}]_{\times}\mathbf{R}_{j}^{\top}\mathbf{R}_{j}[\mathbf{n}]_{\times}^{\top},\\ &=[\mathbf{n}]_{\times}[\mathbf{n}]_{\times}^{\top}.\\ \end{split}$ (12)
@@
$\begin{split}[\mathbf{n}]_{\times}\mathbf{Q}_{j}[\mathbf{n}]_{\times}^{\top}=s_{j}[\mathbf{n}]_{\times}[\mathbf{n}]_{\times}^{\top},\ j=2,3\\ \end{split}$ (16)
@@
To simplify our 12 equations, we can eliminate some unknowns from these equations using the elimination ideal technique [11]. This technique was recently used to solve several minimal camera geometry problems [29].
In our case, we first create an ideal $J$ generated by 12 polynomials (16). Then, the unknown elements of the normal $\{n_{x},n_{y}\}$ and the scale factor $s_{2},s_{3}$ are eliminated from the generators of $J$ by computing the generators of the elimination ideal $J_{1}=J\cap\mathbb{C}[q_{21},\dots,q_{36}]$. Here, $q_{j.}$ are the entries of $\mathbf{Q}_{2},\mathbf{Q}_{3}$. These generators can be computed using computer algebra software Macaulay2 [17] (for more details and the input Macaulay2 code see the SM).
@@ 5 New Solvers
In this section, we propose solvers for the four different cases outlined in Tab. 1 using the constraints derived in the previous section. We also propose a method based on LO-RANSAC [8] that utilizes the new solvers for robust estimation of focal lengths from three views of planar scenes.
@@ 5.1 One Unknown Focal Length Parameter
Case I. We first consider the case where the three cameras have equal and unknown focal length, i.e., $f_{1}=f_{2}=f_{3}=f$, and $\mathbf{K}_{1,2,3}={\rm diag}(f,f,1)$. By substituting $\mathbf{K}_{j}$ into the generators $g_{i}$, $i=1,\dots,7$ of the elimination ideal $J$, we obtain 7 univariate polynomials in $f$, which are of degree 9 in $\alpha=f^{2}$. They form an over-constrained system. We only need one of them to find the solutions to $f$. To find the roots of the degree 9 univariate polynomial we use the Sturm sequence method [16]. We denote this solver as $\mathbf{H}_{fff}$.
Case II. In the second case, we assume that $f_{1}$ is known, and $f_{2}=f_{3}=f$ are unknown. This case occurs when a calibrated camera is used to capture the reference image, and the second uncalibrated camera captures two target images. Similar to Case I, we only need to find the roots of a univariate polynomial in $f$, in this case of degree 6, using Sturm sequences [16]. We denote this solver as $\mathbf{H}_{ff}$.
@@ 5.2 Two Unknown Focal Length Parameters
is a vector consisting of the 28 monomials. The system of polynomial equations in (17) can be solved using different algebraic methods [11]. There are also different state-of-the-art approaches for generating efficient algebraic solvers [26, 32, 2, 27, 19]. In this paper, we use the hidden variable technique to derive polynomial eigenvalue solution based on [27].
@@
where ${\mathbf{C}}_{3},{\mathbf{C}}_{2},{\mathbf{C}}_{1},{\mathbf{C}}_{0}$ are $7\times 7$ coefficient matrices containing only numbers. For this problem, the matrix $\mathbf{C}_{3}$ is only rank 4, resulting in four zero eigenvalues. To speed up the solver, we remove these zero eigenvalues from the computations. To do this, we first need to transform the matrices $\mathbf{C}_{i}$ by considering linear combinations of their rows, such that there are three zero rows in the transformed matrix $\mathbf{C}_{3}$. To remove the zero rows in the transformed $\mathbf{C}_{3}$, we use the technique from [15] and consider the transpose of (19)
@@
If we consider (23) as a polynomial eigenvalue problem [1], the solutions to $\gamma$ are the eigenvalues of $21\times 21$ matrix
@@
The three zero columns in ${\mathbf{C}}_{3}^{\top}$ can now be removed together with their corresponding rows to eliminate the zero eigenvalues [27]. In this way, we obtain 18 possible solutions. Once we have solutions to $\alpha$, the remaining unknown $\beta$ can be extracted from the null vector of $\mathbf{C}(\alpha)$ based on (19). We denote this solver as $\mathbf{H}_{f\rho\rho}$.
Note that in the proposed polynomial eigenvalue formulation, we solve a relaxed version of the original problem (11). The original system (11) has 17 solutions, as it can be shown, e.g., using the computer algebra system Macaulay2 [17]. In the polynomial eigenvalue formulation, we have one spurious solution that does not ensure that the elements of $\tilde{\mathbf{v}}$ satisfy ${\tilde{\mathbf{v}}}=[1,\beta,\cdots,\beta^{6}]^{\top}$.

### v2

@@ 1 Introduction
Estimating relative camera motion from multiple views using point correspondences is a classical problem in computer vision. Efficient solutions for various camera configurations have been well-studied in the literature [nister2004efficient, hartley2012efficient, kukelova2012polynomial, stewenius2005minimal, li2006simple, brown2007minimal, hartley1992estimation, hartley2003multiple, hartley1997defense, nister2006Four]. For example, with two fully calibrated cameras, the relative camera pose can be efficiently determined using 5-point algorithms [nister2004efficient, hartley2012efficient, kukelova2012polynomial]. In contrast to the two view case, the calibrated three-view relative pose problem is much more challenging. This problem can be solved using four triplets of point correspondences [nister2006Four, quan2006Results], [^1] Note, that configuration of four points in three views generates an over-constrained problem. In this case, we have 12 constraint for 11 degrees of freedom (DOF). A minimal solution would need to drop one constraint, e.g., by considering only a line passing through one of the points in the third view [hruby2022Learning] or by considering a “half” point correspondence. however it results in a very complex system of polynomial equations. Thus, the existing solutions [nister2006Four, hruby2022Learning] to this problem are only approximate and can often fail, i.e., the returned solution can be arbitrarily far from the geometrically correct solution.
@@
An important scenario arises when the focal length is the only unknown in the camera intrinsic parameters, referred to as the partially calibrated case. This setup is practical since most modern cameras have zero skew and a centered principal point. When two cameras share an equal yet unknown focal length, their relative motion and the common focal length can be estimated using six point correspondences [stewenius2005minimal, hartley2012efficient, kukelova2012polynomial, kukelova2017clever]. Similarly, if one camera is fully calibrated and the other has an unknown focal length, six point correspondences are also required [bujnak20093d, kukelova2012polynomial, kukelova2017clever]. For two cameras with different, unknown focal lengths, at least seven point correspondences are necessary to recover both the relative motion and the focal lengths [hartley1992estimation, bougnoux1998projective].
Three-view focal length problems are significantly more complex, often involving polynomial systems with hundreds of solutions. For example, when the three cameras share the same focal length, there can be up to 668 possible solutions [chien2022gpu, ding2023minimal, cin2024minimal]. Such systems can be solved using homotopy continuation methods [sommese2005numerical, hauenstein2018adaptive]. The running times of the CPU variants of the solvers for four points in three cameras with unknown shared focal length range from $250ms$ to $1456ms$. Efficient GPU implementations are much faster; however, with runtimes $16.7ms$ to $154ms$ they are still too slow for practical applications.
Scenes with planar surfaces, such as floors, walls, doors, facades, and other common structures, are prevalent in man-made environments. When points are coplanar, homography-based algorithms require four point correspondences [hartley2003multiple] to estimate the relative pose of two cameras. Unfortunately, any attempt to recover intrinsic parameters from two views of a planar surface (using only point correspondences without additional priors) is futile, as stated in the following theorem [maybank1993theory, nister2004efficient]: ”For any choice of intrinsic parameters, any homography can be realized between two views by positioning the views and a plane in a particular way.” However, it has been shown that focal lengths can be recovered from three-view homographies [malis2002camera, heikkila2017using], i.e. three cameras observing a planar surface (see illustration in Figure 1).
@@
  • This allows us to develop solvers using efficient Sturm sequences or hidden variable technique. The proposed solvers are significantly more efficient than the existing solvers to this problem [heikkila2017using]. Moreover, in extensive synthetic and real experiments, we show an improvement in accuracy over the state-of-the-art two-view solvers.
  • We propose a new dataset consisting of 6 scenes (four indoor and 2 outdoor) containing 1870 images captured with 14 different cameras. Ground truth focal lengths are estimated using the standard calibration method [zhang2000calibration]. We will make both the dataset and the code publicly available.
@@ 2 Related Work
Homography estimation is a well-studied problem in the literature and can be solved using the 4-point algorithm, which involves solving a system of eight homogeneous linear equations [hartley2003multiple]. However, with only point correspondences, two-view homography does not provide additional constraints on the focal length, making the two-view homography not suitable for self-calibration. To recover the intrinsic parameters from the two-view homography, we need additional priors. In [brown2007minimal], under pure rotation assumption, the authors propose two minimal solutions for the two-view homography-based focal length estimation with one or two focal length parameters. In [ding2020homography], the authors use sensor-fusion, i.e., combining camera with gravity from IMU (Inertial measurement unit), to reduce the DOF of the motion parameters. In general, the gravity prior can provide up to two constraints on the focal length.
Focal length recovery from homographies in more views was first discussed in [malis2002camera]. Malis et al. compute homographies among multiple views, iteratively testing various focal lengths to find the one that minimizes a cost function. However, this method is time-consuming and sensitive to noise.
The work most closely related to ours is by Heikkilä [heikkila2017using], where the constraints on the focal length and the normal vector are derived, and non-iterative solutions to the problem of recovering focal lengths from three-view homographies is proposed. For three cameras with the same unknown focal length,  [heikkila2017using] solves three polynomials in three unknowns. However, these polynomials have very high degrees, making them challenging to solve efficiently. Heikkilä first applies SVD to a $82\times 82$ matrix, followed by Gaussian-Jordan elimination of a $82\times 164$ matrix. Finally, the solver finds the solutions by computing the eigenvalues of a $82\times 82$ matrix. For three cameras with two different focal lengths, Heikkilä’s method needs to solve four polynomials in four unknowns by performing Gaussian-Jordan elimination of a $176\times 352$ matrix and computing the eigenvalues of a $176\times 176$ matrix.
In contrast to complex solutions presented in [heikkila2017using], we propose very efficient solutions to the problem of recovering focal lengths from three-view homographies. For cameras with the same focal length, our novel solver only requires computing the roots of a univariate polynomial of degree 9, which can be efficiently solved using Sturm sequences. For cameras with two different focal lengths, we compute solutions from the eigenvalues of a $18\times 18$ matrix.
@@
Problems Proposed Solvers View 1 View 2 View 3 Ours Heikkilä [heikkila2017using] Ours Heikkilä [heikkila2017using] Ours Heikkilä [heikkila2017using]
@@
Case IV $\mathbf{H}_{f\rho}$ Known $f$ $\rho$ 9 - $12\times 12$ - 106 -
@@
For general scenes, focal lenghts can be recovered together with the epipolar geometry. For two cameras with equal unknown focal lengths, the minimal 6-point solvers [stewenius2005minimal, bujnak20093d, hartley2012efficient, kukelova2012polynomial, kukelova2017clever] use rank-2 and trace constraints on the essential matrix. The 5-point solver [torii2011six] solves the plane+parallax scenario for cameras with unknown focal length using four coplanar points and one off-plane point. This solver can be used within the DEGENSAC framework [chum2005degensac] to complement the six point solvers in order to improve performance for scenes that are not completely planar, but contain a dominant plane. For the single unknown focal length problem, the two-view geometry can be solved using the single-side 6-point solver [bujnak20093d, kukelova2017clever]. For the relative pose problem with different and unknown focal lengths, the solution computes the fundamental matrix using the 7-point solver [hartley2003multiple], followed by the extraction of the focal lengths from this matrix [bougnoux1998projective, kocur2024robust].
As discussed above, three-view focal length problems for general scenes are significantly more complex and their homotopy continuation solutions [chien2022gpu, ding2023minimal, cin2024minimal] are impractical.
@@ 3 Problem Statement
The points are observed by three cameras such that $\mathbf{X}_{i}$ is projected to a 2D point $\mathbf{m}_{i,j}$ by the $j$-th camera $\mathbf{K}_{j}[\mathbf{R}_{j}\ |\ \mathbf{t}_{j}]$. In many practical scenarios, it is often reasonable to assume that the cameras have square-shaped pixels, and the principal point coincides with the image center [hartley2012efficient]. This leaves only the focal point unknown and thus $\mathbf{K}_{j}={\rm diag}(f_{j},f_{j},1)$.
@@ 4 Our Approach
[…]
Thus for known $\mathbf{G}_{j}$, $\mathbf{H}_{j}$ are polynomial matrices in the focal length parameters. To solve for the focal lengths, we need to find the constraints on $\mathbf{H}_{j}$. Based on the formulation of the Euclidean homography (4), we know that [hartley2003multiple]
@@
Thus the essential matrices derived from different $\mathbf{H}_{j}^{\top}$ are related by the same normal vector $\mathbf{n}$ in the reference camera coordinate. As shown in [stefanovic1973relative, nister2004efficient], a valid essential matrix should satisfy the singular and trace constraints
@@
Substituting (7) into the trace constraints (11) gives us nine equations per homography, i.e., 18 equations for two homography matrices $\mathbf{H}_{2}$ and $\mathbf{H}_{3}$. Only two from the nine equations are algebraically independent, i.e., one homography matrix provides two constraints on the focal lengths and the normal vector parameters. For the case of equal unknown focal length, we have only 3 unknowns $(f,n_{x},n_{y})$ (since (11) is homogenous in $\mathbf{n}$, we can let $n_{z}=1$) and thus an over-constrained problem. The problem can be solved using only one of the nine trace constraint equations for $\tilde{\mathbf{E}}_{2}$. These equations can be solved using the Gröbner basis method [larsson2017efficient, larsson2018beyond] The final solver performs Gauss-Jordan elimination of a $118\times 141$ matrix and extracts solutions from the eigenvectors of a $23\times 23$ matrix. However, this solver is not efficient enough for real-time applications.
@@
[…]
@@
[…]
@@
To simplify our 12 equations, we can eliminate some unknowns from these equations using the elimination ideal technique [cox2005using]. This technique was recently used to solve several minimal camera geometry problems [kukelova2017clever].
In our case, we first create an ideal $J$ generated by 12 polynomials (16). Then, the unknown elements of the normal $\{n_{x},n_{y}\}$ and the scale factor $s_{2},s_{3}$ are eliminated from the generators of $J$ by computing the generators of the elimination ideal $J_{1}=J\cap\mathbb{C}[q_{21},\dots,q_{36}]$. Here, $q_{j.}$ are the entries of $\mathbf{Q}_{2},\mathbf{Q}_{3}$. These generators can be computed using computer algebra software Macaulay2 [M2] (for more details and the input Macaulay2 code see the SM).
@@ 5 New Solvers
In this section, we propose solvers for the four different cases outlined in Tab. 1 using the constraints derived in the previous section. We also propose a method based on LO-RANSAC [chum2003loransac] that utilizes the new solvers for robust estimation of focal lengths from three views of planar scenes.
@@ 5.1 One Unknown Focal Length Parameter
Case I. We first consider the case where the three cameras have equal and unknown focal length, i.e., $f_{1}=f_{2}=f_{3}=f$, and $\mathbf{K}_{1,2,3}={\rm diag}(f,f,1)$. By substituting $\mathbf{K}_{j}$ into the generators $g_{i}$, $i=1,\dots,7$ of the elimination ideal $J$, we obtain 7 univariate polynomials in $f$, which are of degree 9 in $\alpha=f^{2}$. They form an over-constrained system. We only need one of them to find the solutions to $f$. To find the roots of the degree 9 univariate polynomial we use the Sturm sequence method [gellert2012vnr]. We denote this solver as $\mathbf{H}_{fff}$.
Case II. In the second case, we assume that $f_{1}$ is known, and $f_{2}=f_{3}=f$ are unknown. This case occurs when a calibrated camera is used to capture the reference image, and the second uncalibrated camera captures two target images. Similar to Case I, we only need to find the roots of a univariate polynomial in $f$, in this case of degree 6, using Sturm sequences [gellert2012vnr]. We denote this solver as $\mathbf{H}_{ff}$.
@@ 5.2 Two Unknown Focal Length Parameters
is a vector consisting of the 28 monomials. The system of polynomial equations in (17) can be solved using different algebraic methods [cox2005using]. There are also different state-of-the-art approaches for generating efficient algebraic solvers [kukelova2008automatic, larsson2018beyond, Bhayani_2020_CVPR, kukelova2012polynomial, hartley2012efficient]. In this paper, we use the hidden variable technique to derive polynomial eigenvalue solution based on [kukelova2012polynomial].
@@
where ${\mathbf{C}}_{3},{\mathbf{C}}_{2},{\mathbf{C}}_{1},{\mathbf{C}}_{0}$ are $7\times 7$ coefficient matrices containing only numbers. For this problem, the matrix $\mathbf{C}_{3}$ is only rank 4, resulting in four zero eigenvalues. To speed up the solver, we remove these zero eigenvalues from the computations. To do this, we first need to transform the matrices $\mathbf{C}_{i}$ by considering linear combinations of their rows, such that there are three zero rows in the transformed matrix $\mathbf{C}_{3}$. To remove the zero rows in the transformed $\mathbf{C}_{3}$, we use the technique from [ding2020homography] and consider the transpose of (19)
@@
If we consider (23) as a polynomial eigenvalue problem [bai2000templates], the solutions to $\gamma$ are the eigenvalues of $21\times 21$ matrix
@@
The three zero columns in ${\mathbf{C}}_{3}^{\top}$ can now be removed together with their corresponding rows to eliminate the zero eigenvalues [kukelova2012polynomial]. In this way, we obtain 18 possible solutions. Once we have solutions to $\alpha$, the remaining unknown $\beta$ can be extracted from the null vector of $\mathbf{C}(\alpha)$ based on (19). We denote this solver as $\mathbf{H}_{f\rho\rho}$.
Note that in the proposed polynomial eigenvalue formulation, we solve a relaxed version of the original problem (11). The original system (11) has 17 solutions, as it can be shown, e.g., using the computer algebra system Macaulay2 [M2]. In the polynomial eigenvalue formulation, we have one spurious solution that does not ensure that the elements of $\tilde{\mathbf{v}}$ satisfy ${\tilde{\mathbf{v}}}=[1,\beta,\cdots,\beta^{6}]^{\top}$.